# A dynamical approach to studying the Lee-Yang zeros for the Potts Model on the Cayley Tree

### v1

@@ 3. PROOF OF THEOREM A
then the orbit $\{f^{n}(z_{0})\}_{n=0}^{\infty}$ of $x_{0}$ converges to $x_{N}$.
@@
Let $t\in[0,\infty)\setminus\{1\}$ and $z\geq 0$. Then the marked point $a(z)=z$ is a fixed point of $R_{z,t,q}(w)$ if and only if $z=1$.
@@
$Q(z,t,q):=$ $\left(4q^{3}t^{4}-24q^{2}t^{4}+48qt^{4}-24(q-2)t^{3}+12(q-2)qt^{3}+12qt^{2}-32t^{4}-24t^{2}+4t\right)z^{2}$
$\left(-q^{2}t^{4}+3q^{2}t^{2}+2qt^{4}-6(q-2)t^{3}+6(q-2)qt^{3}-6qt^{2}+6qt-t^{4}+6t^{2}-12t+3\right)z$
$4qt-4t=0.$
Since $t\in[0,1)$, $z>0$, and $q\geq 2$ the only way that the discriminant of $P_{z,t,q}(w)$ vanishes is when $Q(z,t,q)$ vanishes. The reader can check that $Q(z,t,q)$ vanishes precisely under the conditions stated in the conclusion of this lemma.
@@
Suppose $t_{1}(q)<t<t_{2}(q)$ and $z=N_{-}(t,q)$. By Lemma 12(ii(b)), the orbit of $z_{0}=N_{-}(t,q)$ converges to the neutral fixed point $w_{N}(N_{-}(t,q))$. Furthermore, there exists an attracting fixed point $w_{A}<w_{N}(N_{-}(t,q))$. This attracting fixed point varies continuously with respect to $z$ around $N_{-}(t,q)$. We will denote this dependence on $z$ by $w_{A}\equiv w_{A}(z)$. By Lemma 12(iii) if $z<z_{0}$ is sufficiently close to $z_{0}$, then the orbit of $z$ converges to $w_{A}(z)$. Thus, for any subsequence $\left(R_{z,t,q}^{n_{k}}(z)\right)_{k=1}^{\infty}$ of $\left(R_{z,t,q}^{n}(z)\right)_{n=1}^{\infty}$ and any neighborhood $U$ of $z_{0}=N_{-}(t,q)$, we can find a further subsequence $\left(R_{z,t,q}^{n_{k_{j}}}(z)\right)_{j=1}^{\infty}$ of $\left(R_{z,t,q}^{n_{k}}(z)\right)_{k=1}^{\infty}$ and a point $z_{1}\in U$ such that for all $j\in\mathbb{N}$
@@ 4. Proof of Theorem B (Antiferromagnetic Case, $q\geq 3$)
(15) $\mathcal{A}(R_{z,t,q})\cap(0,\infty)=\begin{cases}\emptyset\ &\mbox{if }\ 1\leq t<t_{3}(q)\\ \{z_{c}^{\pm}(t,q)\}\ &\mbox{if }\ t_{3}(q)\leq t.\end{cases}$
@@
To prove additional claim (b) it suffices to show that $Q_{1}(z,t,q)>0$ for all $z\geq 1$, $t\geq 1$, and $q\geq 3$. If one computes $Q_{1}(1+r,1+s,3+p)$ one finds a polynomial whose constant term is $72$ and all of whose monomials appear with a “plus” sign. (We have omitted the many line expression which is similar to Equation (4) above, but the reader can check it in a computer algebra software package.). ∎
@@
$w_{1}$ $:=\frac{939-17\sqrt{2165}}{1058}\approx 0.1399,\qquad w_{2}:=\frac{939+17\sqrt{2165}}{1058}\approx 1.6352\quad\mbox{and}$
@@ 5. Proof of Theorem C (Antiferromagnetic Case, $q=2$)
(20) $\mathcal{A}(R_{z,t,2})\cap(0,\infty)=\begin{cases}\emptyset\ &\mbox{if }\ 1\leq t<3\\ \{1\}\cup\{z_{c}^{\pm}(t,2)\}\ &\mbox{if }\ 3\leq t.\end{cases}$
@@
Proof of Claim (iii): First suppose that $z=N_{-}(t,q)$ for $(t,q)$ in Region II, as discussed in Case (ii) above. Since the graph of $R_{z,t,q}(w)$ does not cross the diagonal at the fixed point $w_{N}(t,q)$ we have that for all $w\in(w_{A}(t,q),\infty)$ that $R_{z,t,q}(w)\leq w$ with equality occurring at $w=w_{N}(t,q)$. Decreasing the parameter $z>0$ decreases the values of $R_{z,t,q}(w)$ and this will eliminate the neutral fixed point $w_{N}(t,q)$. Meanwhile, if the decrease of the parameter $z>0$ is by a sufficiently small amount, then the then the attracting fixed point $w_{A}(t,q)$ will move continuously to an attracting fixed point $w_{A}^{\prime}$ for the perturbed map, and we will have that $R_{z,t,q}(w)<w$ for all $w\in(w_{A}^{\prime},\infty)$. This implies that the orbit of the marked point $w=z$ will now converge to the new attracting fixed point $w_{A}^{\prime}$, as claimed in Part (iii) of the Lemma. ∎
@@ Appendix B Derivation of the Renormalization Mapping (Proof of Theorem D)
Here $W_{n}(\sigma):=e^{-\frac{H_{n}(\sigma)}{T}}$ is the Boltzmann-Gibbs weight of the configuration $\sigma$.
(21) Notice that $Z_{n}^{j}=Z_{n}^{k}$ for any $1\leq j,k\leq q-1$.

### v2

@@ 3. PROOF OF THEOREM A
then the orbit $\{f^{n}(x_{0})\}_{n=0}^{\infty}$ of $x_{0}$ converges to $x_{N}$.
@@
Let $t\in(0,\infty)\setminus\{1\}$ and $z>0$. Then the marked point $a(z)=z$ is a fixed point of $R_{z,t,q}(w)$ if and only if $z=1$.
@@
$Q(z,t,q):=$ $\left(4q^{3}t^{4}-24q^{2}t^{4}+48qt^{4}-24(q-2)t^{3}+12(q-2)qt^{3}+12qt^{2}-32t^{4}-24t^{2}+4t\right)z^{2}+$
$\left(-q^{2}t^{4}+3q^{2}t^{2}+2qt^{4}-6(q-2)t^{3}+6(q-2)qt^{3}-6qt^{2}+6qt-t^{4}+6t^{2}-12t+3\right)z+$
$4qt-4t.$
Since $t\in(0,1)$, $z>0$, and $q\geq 2$ the only way that the discriminant of $P_{z,t,q}(w)$ vanishes is when $Q(z,t,q)$ vanishes. The reader can check that $Q(z,t,q)$ vanishes precisely under the conditions stated in the conclusion of this lemma.
@@
Suppose $t_{1}(q)<t<t_{2}(q)$ and $z=N_{-}(t,q)$. By Lemma 12(ii(b)), the orbit of $z_{0}=N_{-}(t,q)$ converges to the neutral fixed point $w_{N}(N_{-}(t,q))$. Furthermore, there exists an attracting fixed point $0<w_{A}<w_{N}(N_{-}(t,q))$. This attracting fixed point varies continuously with respect to $z$ around $N_{-}(t,q)$. We will denote this dependence on $z$ by $w_{A}\equiv w_{A}(z)$. By Lemma 12(iii) if $z<z_{0}$ is sufficiently close to $z_{0}$, then the orbit of $z$ converges to $w_{A}(z)$. Thus, for any subsequence $\left(R_{z,t,q}^{n_{k}}(z)\right)_{k=1}^{\infty}$ of $\left(R_{z,t,q}^{n}(z)\right)_{n=1}^{\infty}$ and any neighborhood $U$ of $z_{0}=N_{-}(t,q)$, we can find a further subsequence $\left(R_{z,t,q}^{n_{k_{j}}}(z)\right)_{j=1}^{\infty}$ of $\left(R_{z,t,q}^{n_{k}}(z)\right)_{k=1}^{\infty}$ and a point $z_{1}\in U$ such that for all $j\in\mathbb{N}$
@@ 4. Proof of Theorem B (Antiferromagnetic Case, $q\geq 3$)
(15) $\mathcal{A}(R_{z,t,q})\cap(0,\infty)=\begin{cases}\emptyset\ &\mbox{if }\ 1<t<t_{3}(q)\\ \{z_{c}^{\pm}(t,q)\}\ &\mbox{if }\ t_{3}(q)\leq t.\end{cases}$
@@
To prove additional claim (b) it suffices to show that $Q_{1}(z,t,q)>0$ for all $z\geq 1$, $t\geq 1$, and $q\geq 3$. If one computes $Q_{1}(1+r,1+s,3+p)$ one finds a polynomial whose constant term is $72$ and all of whose monomials appear with a “plus” sign. (We have omitted the many line expression which is similar to Equation (4) above, but the reader can check it in a computer algebra software package.) ∎
@@
$w_{1}$ $:=\frac{939-17\sqrt{2165}}{1058}\approx 0.1399,\qquad w_{2}:=\frac{939+17\sqrt{2165}}{1058}\approx 1.6352,\quad\mbox{and}$
@@ 5. Proof of Theorem C (Antiferromagnetic Case, $q=2$)
(20) $\mathcal{A}(R_{z,t,2})\cap(0,\infty)=\begin{cases}\emptyset\ &\mbox{if }\ 1<t<3\\ \{1\}\cup\{z_{c}^{\pm}(t,2)\}\ &\mbox{if }\ 3\leq t.\end{cases}$
@@
Proof of Claim (iii): First suppose that $z=N_{-}(t,q)$ for $(t,q)$ in Region II, as discussed in Case (ii) above. Since the graph of $R_{z,t,q}(w)$ does not cross the diagonal at the fixed point $w_{N}(t,q)$ we have that for all $w\in(w_{A}(t,q),\infty)$ that $R_{z,t,q}(w)\leq w$ with equality occurring at $w=w_{N}(t,q)$. Decreasing the parameter $z>0$ decreases the values of $R_{z,t,q}(w)$ and this will eliminate the neutral fixed point $w_{N}(t,q)$. Meanwhile, if the decrease of the parameter $z>0$ is by a sufficiently small amount, then the attracting fixed point $w_{A}(t,q)$ will move continuously to an attracting fixed point $w_{A}^{\prime}$ for the perturbed map, and we will have that $R_{z,t,q}(w)<w$ for all $w\in(w_{A}^{\prime},\infty)$. This implies that the orbit of the marked point $w=z$ will now converge to the new attracting fixed point $w_{A}^{\prime}$, as claimed in Part (iii) of the Lemma. ∎
@@ Appendix B Derivation of the Renormalization Mapping (Proof of Theorem D)
Here $W_{n}(\sigma):=e^{-\frac{H_{n}(\sigma)}{T}}$ is the Boltzmann-Gibbs weight of the configuration $\sigma$. Notice that
(21) $Z_{n}^{j}=Z_{n}^{k}$ for any $1\leq j,k\leq q-1$.
@@
Disclosures and declarations:
As required by current rules for articles submitted to Springer-Nature journals, the authors declare that:
  1. (1)
    The authors have no financial or non-financial or competing interests that are relevant to this article. The authors have stated their grant support in the Acknowledgments.
  2. (2)
    Data accessibility: this article is theoretical and does not contain any experimental data. Calculational data supporting the conclusions of the article are included in the text of the article.